OpenStudy (anonymous):

A problem in my testbok says to rewrite the expression 1/4 of 1/4 of 32 and the answer is 1/4*1/4*32=32/16=2. Could sombody explain to me how the answer is true?

OpenStudy (anonymous):

"1/4 of" something means "1/4" times that something So, 1/4 of 1/4 of 32 = (1/4)*(1/4)*32 = (1/16)*32 = 2

OpenStudy (anonymous):

Firstly, lets go to the second part of the question. 1/4 of 32 is 1/4*32 = 8. From there, we work at the first part which is "1/4 of" 1/4 of 32. 1/4 of 8 is 2. Hence, the ans.

OpenStudy (anonymous):

Oh, right. Ok thanks.

OpenStudy (anonymous):